What is the change in Gibb's Energy, in kJ/mol, of a melting ice cube at 20 degrees Celsius? At -10 degrees celsius? Please show work.

The answer was -0.44 for at 20 degrees Celsius. Not sure how to get there.
Answer #1

Ans:-H2O(s)-----> H20(l).

melting of ice at 20°c

∆Hfus= 6.01 KJ/mol

∆S= 22.0 J/mol

T = 20 + 273= 293K

∆G= ∆H-T∆S

∆G= 6.01- 293 × 0.022 KJ/mol

∆G = -0.44 KJ/mol.

∆G <0 means spontaneous process

Melting of ice at -10°c

T= 263K

∆G= ∆H -T∆S

∆G = 6.01- 263×0.022 KJ/mol

∆G= 0.22 KJ/mol

∆G > 0 means not spontaneous rxn will reverse .